OpenStudy (anonymous):

what is 4/10+3/10+2/10-1/9

OpenStudy (anonymous):

Do you know how to find a common denominator?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Do you know how to add fractions of the same denominator?

OpenStudy (anonymous):

Well an example of making a common denominator is this. 1/2 + 1/3. You can't add them together yet, so you look at what they're equal to. 1/2 = 2/4 = 3/6 and so on 1/3 = 2/6 = 3/9 and so on. When you compare the two lists, the first in each that have the same denominator, in this case 3/6 and 2/6, you use them in your problem.

OpenStudy (anonymous):

to find a common denominator without worrying about the smallest one, you can just multiply like this. 1/2 * 3/3 + 1/3*2/2. You can do this because you're just multiplying by 1. This will also put both numbers over 6.